OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (−1, −13); perpendicular to the line passing through (2, −1) and (6, −3)

OpenStudy (owlcoffee):

Okay so, we have a point and a line that goes through those two points. Write the ecuation of the first line and then deduce the slope for both, use the ecuation for a line intersecting a point and you're done.

OpenStudy (anonymous):

can you help

OpenStudy (owlcoffee):

Sure. Okay so, let's first look at the line that intersects two points. we'll determine it's slope: \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] Can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

?

OpenStudy (owlcoffee):

The ecuation would have this form: \[m=\frac{ (-3)-(-1) }{ 6-2 }\] Giving me a result of : m= -1/2 Now, using that slope, we should deduce the slope of the second line with this: \[m _{1}=-\frac{ 1 }{ m _{2} }\] That means that the other slope, should be opposite to the one we just found, what value would that be for M2?

OpenStudy (anonymous):

2

OpenStudy (owlcoffee):

Great, now, the problem gives us the point, the line having that slope intersects, So let's use this ecuation: \[y-y _{1}=m(x-x _{1})\] Where the x1 and y1 terms are the points we want to find.

OpenStudy (anonymous):

how do i do that

OpenStudy (owlcoffee):

the problem said the line whent through (-1,-13) So the ecuation would look like this: \[y+13=2(x+1)\]

OpenStudy (anonymous):

so it would be y=2x+15

OpenStudy (owlcoffee):

Correct!

OpenStudy (anonymous):

Some scientists believe that the average surface temperature of the world has been rising steadily. The average surface temperature can be modeled by T = 0.02t + 15.0 Use the equation to predict the average global surface temperature in 2090. (Round your answer to the nearest degree.)

OpenStudy (owlcoffee):

Well, ther you have to simply replace